Find the difference quotient for -f(x)= x^2-3x+5
Answer by stanbon(75887)

Find the difference quotient for -f(x)= x^2-3x+5
--------------
f(x)= -x^2+3x-5
f(x+h) = -[x^2+2hx+h^2]+3(x+h)-5
f(x+h) = -x^2-2hx-h^2+3x+3h-5
---------------
f(x+h)-f(x) = -2hx-h^2+3h
-----------------
Difference quotient = [f(x+h)-f(x)]/h
---------------
[f(x+h)-f(x)]/h = -2x-h+3
===================
Cheers,
Stan H.